# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 3. Local well-posedness
In this section, we prove our main local well-posedness result, Theorem 1.1. Throughout this section, we fix $\alpha,\beta\in\mathbb{R}$ and do not indicate the dependence of constants on these parameters.
We employ a compactness argument. For $\ell\in\mathbb{Z}_{\geq 0}$, consider the mollified equation for (gINLS):
@@
The proof proceeds in four main steps. First, in Proposition 3.1, we establish the existence of a unique local solution $u^{(\ell)}\in C([0,T_{\ell});\mathcal{Z}^{\infty}_{\rho})$ to the mollified problem (3.1) with mollified initial condition $u^{(\ell)}(0)=P_{\leq\ell}\phi$. Second, in Proposition 3.2, we derive uniform-in-$\ell$ existence time $T_{0}>0$ and $\mathcal{Z}^{2}_{\rho}$ a priori bound for the sequence $\{u^{(\ell)}\}_{\ell\geq 0}$, valid for all $t\in[0,T_{0}]$. Third, in Proposition 3.3, we prove that $\{u^{(\ell)}\}_{\ell\geq 0}$ is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$ for some possibly shorter time $T_{*}\leq T_{0}$. Finally, we pass to the limit $\ell\to\infty$ and prove uniqueness and continuity of the flow map in $\mathcal{Z}^{2}_{\rho}$, completing the proof of Theorem 1.1.
@@
Fix $\rho>0$, and let $\phi\in\mathcal{Z}^{1}_{\rho}$. For any $\ell\in\mathbb{Z}_{\geq 0}$, the initial value problem (3.1) with initial condition $u(0)=P_{\leq\ell}\phi$ admits a unique maximal-lifespan solution $u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}^{1}_{\rho})$ which belongs to $C([0,T_{\ell}(\phi));\mathcal{Z}^{\infty}_{\rho})$. Moreover, if $T_{\ell}(\phi)<\infty$, then $\lim\limits_{t\to T_{\ell}(\phi)}\mathcal{E}^{1}_{\rho}(u^{(\ell)}(t))=\infty$.
Remark 3.1.
From the proof below and the estimate $\mathcal{E}^{1}_{\rho}(P_{\leq\ell}\phi)\leq 3\mathcal{E}^{1}_{\rho}(\phi)$ from (2.9), we see that $T_{\ell}(\phi)\gtrsim\big[(\delta^{-1}+2^{\ell})(1+[\mathcal{E}^{1}_{\rho}(\phi)]^{2})\big]^{-1}$. This does not imply a uniform-in-$\ell$ lower bound, however.
@@
It suffices to construct a unique local solution $u^{(\ell)}\in C([0,T];\mathcal{Z}^{1}_{\rho})$ to the initial value problem of (3.1) with initial data $\phi$ satisfying $\mathrm{supp}\,(\widehat{\phi})\subset\{|\xi|\leq 2^{\ell+1}\}$ and $\mathcal{E}^{1}_{\rho}(\phi)\leq M_{0}^{2}$, for some $T>0$ depending on $M_{0}$ and $\ell$. Indeed, the solution $u^{(\ell)}$ then satisfies $\widehat{u^{(\ell)}}(\xi,t)=0$ outside $\{|\xi|\leq 2^{\ell+1}\}$ for all $t\geq 0$, from which the regularity of $u^{(\ell)}$ follows.
Recall from (2.7) and (2.19) that $\|\mathcal{E}^{1}_{\rho}(U(t)\phi+\Phi^{(\ell)}(u))\|_{L^{\infty}(0,1)}\leq C_{*}\mathcal{E}^{1}_{\rho}(\phi)+C\|\Phi^{(\ell)}(u)\|_{L^{\infty}(0,1;H^{1})}^{2}$ for some $C_{*},C>0$. For $0<T\leq 1$, define the complete metric space $(X_{T},d_{T})$ by
$X_{T}:=\big\{u\in C([0,T];\mathcal{Z}^{1}_{\rho}):\|\mathcal{E}^{1}_{\rho}(u)\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}\big\},\qquad d_{T}(u,v):=\|d^{1}_{\rho}(u,v)\|_{L^{\infty}_{T}},$
@@
By (2.7) and (3.3), we have $\|\mathcal{E}^{1}_{\rho}(\mathcal{S}^{(\ell)}_{\phi}[u])\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}$ if $T>0$ is chosen as $T(\delta^{-1}+2^{\ell})M_{0}^{2}\ll 1$. Moreover, using (2.8), (2.20) and (2.16), we can show that the mapping $t\mapsto\mathcal{S}^{(\ell)}_{\phi}[u](t)$ is continuous in $\mathcal{Z}^{1}_{\rho}$, and hence $\mathcal{S}^{(\ell)}_{\phi}$ maps $X_{T}$ into itself. Furthermore, (2.8) with (2.17), (3.3) and (3.4) imply that
@@
Therefore, $\mathcal{S}^{(\ell)}_{\phi}$ is a contraction on $X_{T}$ if $T$ is chosen as $T(\delta^{-1}+2^{\ell})(1+M_{0}^{4})\ll 1$, and the contraction mapping theorem shows that there exists a unique solution $u^{(\ell)}\in X_{T}$ to the initial value problem. The uniqueness in the whole space $C([0,T];\mathcal{Z}^{1}_{\rho})$ can be shown by a standard argument; we have thus proven the proposition. ∎
Proposition 3.2 (Uniform $\mathcal{Z}^{2}_{\rho}$ estimate).
Let $\rho>0$, $M_{0}>0$, and $\phi\in\mathcal{Z}_{\rho}^{2}$ be such that $\mathcal{E}^{2}_{\rho}(\phi)\leq M_{0}^{2}$. For $\ell\geq 0$, let $u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}_{\rho}^{\infty})$ be the unique solution to (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ given in Proposition 3.1. Then, there exists $T_{0}\gtrsim\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ such that for all $\ell$, we have $T_{\ell}(\phi)>T_{0}$ and
(3.5) $\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim\mathcal{E}^{2}_{\rho}(\phi).$
@@
(3.6) $\|\mathcal{E}^{k}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim_{k}\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}_{\rho}(\phi).$
@@
To establish the uniform bound (3.5) in $\mathcal{Z}^{2}_{\rho}$, the principal difficulty arises from a derivative loss in the highest-order energy estimate. To overcome this, we employ a modified energy method: we design a frequency-localized correction term that neutralizes the problematic nonlinear interactions in the nonlinearity of (3.1). In the following proof, we put
@@
Let $\ell\geq 0$, and take arbitrary $T\in(0,1]$ with $T<T_{\ell}(\phi)$. We begin by controlling the lower-order part $\mathcal{E}^{1}_{\rho}(u^{(\ell)})$ using the integral equation (3.2) and (2.7) with $k=1$,
(3.7) $\|\mathcal{E}^{1}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}$ $\lesssim\|\mathcal{E}^{1}_{\rho}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}H^{1}}^{2}$
$\leq\|\mathcal{E}^{1}_{\rho}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\Big(\int_{0}^{T}\|\mathcal{N}(\zeta_{\ell}(s))\|_{H^{1}}\,ds\Big)^{2}$
$\lesssim\mathcal{E}^{1}_{\rho}(\phi)+T^{2}(1+\delta^{-1})^{2}\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3},$
where we have used estimates in Sections 2.3, 2.4, and Lemma 2.1.
Now, we move on to the energy estimate for the highest-order component. Note that the following formal computations are justified since $u^{(\ell)}$ is a smooth solution. Using the equation (3.1), we find
@@
$\Big|\big({-}i\partial_{x}^{2}\mathcal{N}_{\delta}(\zeta_{\ell}(t)),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|\lesssim\delta^{-1}\big[\mathcal{E}^{2}_{\rho}(u(t))\big]^{2}.$
@@
(3.8) $\frac{d}{dt}\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}+C(1+\delta^{-1})\big[\mathcal{E}^{2}_{\rho}(u(t))\big]^{2}.$
The first term on the right-hand side contains a derivative loss that cannot be handled by the standard energy method, and we introduce a time-dependent corrector to compensate with this term. For a frequency threshold $n_{0}\in\mathbb{Z}_{\geq 0}$ to be chosen sufficiently large later, define
@@
This corrector is designed so that its time derivative cancels the above problematic term in the evolution of $\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}$. Expanding the product in (3.9), we obtain the bound
@@
$\lesssim 2^{-n_{0}}\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}.$
Differentiating $\mathcal{I}_{2}(t)$ and substituting the equation (3.1), and using the symmetry property (2.5) for $H\partial_{x}$, we see that
@@
$\quad-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))}\big),\partial_{x}\big(\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)\big)_{L^{2}},$
where for the first and the last terms we have applied an integration by parts. Using Lemma 2.1, we can control these terms as
$|\mathcal{R}(t)|\lesssim\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{2}+(1+\delta^{-1})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{3}.$
Moreover, the main term in $\partial_{t}\mathcal{I}_{2}(t)$ compensates with the first term in (3.8), that is,
@@
(3.11) $\frac{d}{dt}\mathcal{I}_{2}(t)$ $=-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}$
$\quad+C(1+2^{n_{0}})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{3}.$
@@
$\frac{d}{dt}\big(\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{2}(t)\big)$ $\leq C(1+\delta^{-1}+2^{n_{0}})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{3}.$
Then, integrating it from 0 to $t$, and incorporating the bound (3.10) of $\mathcal{I}_{2}(t)$, we see that
(3.12) $\|\partial_{x}^{2}u^{(\ell)}\|_{L^{\infty}_{T}L^{2}}^{2}\leq\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+C\Big\{\big(2^{-n_{0}}+T(1+\delta^{-1}+2^{n_{0}})\big)\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}+T(1+\delta^{-1})\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3}\Big\}.$
@@
(3.13) $\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}_{\rho}(\phi)+\Big\{2^{-n_{0}}\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}+T(1+\delta^{-1}+2^{n_{0}})\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}$
$+\big[T(1+\delta^{-1})+T^{2}(1+\delta^{-1})^{2}\big]\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}\Big\}$ $\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}.$
We now choose $n_{0}$ as $2^{n_{0}}=C(1+M_{0}^{2})$ with a large constant $C>0$ so that $2^{-n_{0}}\mathcal{E}^{2}_{\rho}(\phi)\leq 2^{-n_{0}}M_{0}^{2}\ll 1$, and choose $T_{0}$ sufficiently small so that $T_{0}(1+\delta^{-1})\big((1+2^{n_{0}})M_{0}^{2}+M_{0}^{4}\big)\ll 1$. Then, (3.13) and a standard continuity argument shows that
$\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}_{\rho}(\phi),\qquad 0<T<\min\{T_{0},T_{\ell}(\phi)\}.$
@@
The a priori bound in high regularity (3.6) can be derived by a similar modified energy argument; we only outline the proof. It suffices to estimate the highest component $\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}$. Arguing similarly to the proof of (3.5), by Lemmas 2.1, 2.2 and interpolation inequalities, we deduce that
(3.14) $\frac{d}{dt}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}+C_{k}(1+\delta^{-1})\mathcal{E}^{2}_{\rho}(u(t))\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}.$
@@
$|\mathcal{I}_{k}(t)|\lesssim_{k}2^{-n_{1}}\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},$
$\begin{aligned} \frac{d}{dt}\mathcal{I}_{k}(t)&=-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}\\ &\quad+C_{k}\Big\{(1+2^{n_{1}})\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))+(1+\delta^{-1})\big[\mathcal{E}^{2}_{\rho}(u^{(\ell)}(t))\big]^{2}\Big\}\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}.\end{aligned}$

### v2

@@ 3. Local well-posedness
In this section, we prove our main local well-posedness result, Theorem 1.1. Note that Corollary 1.2 can be proven by the same argument but omitting the nonlinear term $\mathcal{N}_{\delta}(\cdot)$ in the calculations. At the end, we will give a proof of Corollary 1.3. Throughout this section, we fix $\alpha,\beta\in\mathbb{R}$ and do not indicate the dependence of constants on these parameters.
We employ a compactness argument. For $\ell\in\mathbb{Z}_{\geq 0}$, consider the mollified equation for (gINLS) with the mollification parameter $\ell$:
@@
The proof proceeds in four main steps. First, in Proposition 3.1, we establish the existence of a unique local solution $u^{(\ell)}\in C([0,T_{\ell});\mathcal{Z}^{\infty})$ to the mollified problem (3.1) with mollified initial condition $u^{(\ell)}(0)=P_{\leq\ell}\phi$. Second, in Proposition 3.2, we derive uniform-in-$\ell$ existence time $T_{0}>0$ and $\mathcal{Z}^{2}$ a priori bound for the sequence $\{u^{(\ell)}\}_{\ell\geq 0}$, valid for all $t\in[0,T_{0}]$. Third, in Proposition 3.3, we prove that $\{u^{(\ell)}\}_{\ell\geq 0}$ is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1})$ for some possibly shorter time $T_{*}\leq T_{0}$, which gives us a solution of (gINLS) on $0,T_{*}]$ by taking the limit $\ell\to\infty$. Finally, we prove uniqueness and continuity of the flow map in $\mathcal{Z}^{2}$, completing the proof of Theorem 1.1.
@@
Let $\phi\in\mathcal{Z}^{1}$. For each $\ell\in\mathbb{Z}_{\geq 0}$, the Cauchy problem (3.1) with $u(0)=P_{\leq\ell}\phi$ admits a unique maximal solution
$u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}^{\infty}),$
where $T_{\ell}(\phi)$ is the maximal existence time of $u^{(\ell)}$ for which the following blow-up criterion holds: if $T_{\ell}(\phi)<\infty$, then
$\lim_{t\uparrow T_{\ell}(\phi)}\mathcal{E}^{1}(u^{(\ell)}(t))=\infty.$
@@
It is sufficient to construct a unique local solution $u^{(\ell)}\in C([0,T];\mathcal{Z}^{1})$ to the initial value problem of (3.1) with initial data $\phi$ satisfying $\mathrm{supp}\,(\widehat{\phi})\subset\{|\xi|\leq 2^{\ell+1}\}$ and $\mathcal{E}^{1}(\phi)\leq M_{0}^{2}$, for some $T>0$ depending on $M_{0}$ and $\ell$. Indeed, the solution $u^{(\ell)}$ then satisfies $\widehat{u^{(\ell)}}(\xi,t)=0$ outside $\{|\xi|\leq 2^{\ell+1}\}$ for all $t\geq 0$, from which the regularity of $u^{(\ell)}$ follows.
Recall from (2.7) and (2.17) that $\|\mathcal{E}^{1}(U(t)\phi+\Phi^{(\ell)}(u))\|_{L^{\infty}(0,1)}\leq C_{*}\mathcal{E}^{1}(\phi)+C\|\Phi^{(\ell)}(u)\|_{L^{\infty}(0,1;H^{1})}^{2}$ for some $C_{*},C>0$. For $0<T\leq 1$, define the complete metric space $(X_{T},d_{T})$ by
$X_{T}:=\big\{u\in C([0,T];\mathcal{Z}^{1}):\|\mathcal{E}^{1}(u)\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}\big\},\qquad d_{T}(u,v):=\|d^{1}(u,v)\|_{L^{\infty}_{T}},$
@@
By (2.7) and (3.3), we have $\|\mathcal{E}^{1}(\mathcal{S}^{(\ell)}_{\phi}[u])\|_{L^{\infty}_{T}}\leq 2C_{*}M_{0}^{2}$ if $T>0$ is chosen as $T(\delta^{-1}+2^{\ell})M_{0}^{2}\ll 1$. Moreover, using (2.8), (2.18) and (2.14), we can show that the mapping $t\mapsto\mathcal{S}^{(\ell)}_{\phi}[u](t)$ is continuous in $\mathcal{Z}^{1}$, and hence $\mathcal{S}^{(\ell)}_{\phi}$ maps $X_{T}$ into itself. Furthermore, (2.8) with (2.15), (3.3) and (3.4) imply that
@@
Therefore, $\mathcal{S}^{(\ell)}_{\phi}$ is a contraction on $X_{T}$ if $T$ is chosen as $T(\delta^{-1}+2^{\ell})(1+M_{0}^{4})\ll 1$, and the contraction mapping theorem shows that there exists a unique solution $u^{(\ell)}\in X_{T}$ to the initial value problem. The uniqueness in the whole space $C([0,T];\mathcal{Z}^{1})$ and the blow-up criterion can be shown by a standard argument; we have thus proven the proposition. ∎
Proposition 3.2 (Uniform $\mathcal{Z}^{2}$ estimate).
Let $M_{0}>0$, and $\phi\in\mathcal{Z}^{2}$ be such that $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$. For $\ell\geq 0$, let $u^{(\ell)}\in C([0,T_{\ell}(\phi));\mathcal{Z}^{\infty})$ be the unique forward maximal solution to (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ given in Proposition 3.1. Then, there exists $\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}\lesssim T_{0}\leq 1$ such that for all $\ell$, we have $T_{\ell}(\phi)>T_{0}$ and
(3.5) $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim\mathcal{E}^{2}(\phi).$
@@
(3.6) $\|\mathcal{E}^{k}(u^{(\ell)})\|_{L^{\infty}_{T_{0}}}\lesssim_{k}\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}(\phi).$
@@
To establish the uniform bound (3.5), the principal difficulty arises from a derivative loss in the highest-order energy estimate. To overcome this, we employ a modified energy method: we design a frequency-localized correction term that neutralizes the problematic nonlinear interactions in the nonlinearity of (3.1).
Let $\ell\geq 0$, and take arbitrary $T\in(0,1]$ with $T<T_{\ell}(\phi)$. In the following, we put
@@
We begin by controlling the lower-order part $\mathcal{E}^{1}(u^{(\ell)})$ using the integral equation (3.2) and (2.7) with $k=1$,
(3.7) $\|\mathcal{E}^{1}(u^{(\ell)})\|_{L^{\infty}_{T}}$ $\lesssim\|\mathcal{E}^{1}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}H^{1}}^{2}$
$\leq\|\mathcal{E}^{1}(U(t)P_{\leq\ell}\phi)\|_{L^{\infty}_{T}}+\Big(\int_{0}^{T}\|\mathcal{N}(\zeta_{\ell}(s))\|_{H^{1}}\,ds\Big)^{2}$
$\lesssim\mathcal{E}^{1}(\phi)+T^{2}(1+\delta^{-1})^{2}\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3},$
where we have used (2.17), (2.9) and Lemma 2.1 to derive the last inequality.
Now, we move on to the energy estimate for the highest-order component of $\mathcal{E}^{2}(u^{(\ell)})$. Note that the following formal computations are justified since $u^{(\ell)}$ is a smooth solution. Using the equation (3.1), we find
@@
$\Big|\big({-}i\partial_{x}^{2}\mathcal{N}_{\delta}(\zeta_{\ell}(t)),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}\Big|\lesssim\delta^{-1}\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
@@
(3.8) $\frac{d}{dt}\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}.$
The first term on the right-hand side of (3.8) is problematic, as it contains a derivative loss that cannot be handled by the standard energy method. We introduce a time-dependent corrector to compensate with this term. For a frequency threshold $n_{0}\in\mathbb{Z}_{\geq 0}$ to be chosen sufficiently large later, define
@@
This corrector is designed so that its time derivative cancels the problematic term. By Bernstein’s inequality, we obtain the bound
@@
$\lesssim 2^{-n_{0}}\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}.$
Differentiating $\mathcal{I}_{2}(t)$ and using the equation (3.1) and the symmetry property (2.5) for $H\partial_{x}$, we see that
@@
$\quad-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\cdot\partial_{x}\overline{P_{\geq n_{0}}P_{\leq\ell}^{2}\mathcal{N}(\zeta_{\ell}(t))}\big),\partial_{x}\big(\overline{\zeta_{\ell}(t)}\partial_{x}P_{\geq n_{0}}\zeta_{\ell}(t)\big)\big)_{L^{2}}.$
Using Lemma 2.1, we see that
$|\mathcal{R}(t)|\lesssim\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
Moreover, the main term in $\frac{d}{dt}\mathcal{I}_{2}(t)$ compensates with the first term in (3.8), that is,
@@
(3.11) $\frac{d}{dt}\mathcal{I}_{2}(t)$ $\leq-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{3}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{2}\zeta_{\ell}(t)\big)_{L^{2}}$
$\quad+C(1+2^{n_{0}})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
@@
$\frac{d}{dt}\big(\|\partial_{x}^{2}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{2}(t)\big)$ $\leq C(1+\delta^{-1}+2^{n_{0}})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}+C(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{3}.$
Then, integrating it from 0 to $t$, and using the bound (3.10) of $\mathcal{I}_{2}(t)$, we see that
(3.12) $\|\partial_{x}^{2}u^{(\ell)}\|_{L^{\infty}_{T}L^{2}}^{2}\leq\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+C\Big\{\big(2^{-n_{0}}+T(1+\delta^{-1}+2^{n_{0}})\big)\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}+T(1+\delta^{-1})\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{3}\Big\}.$
@@
(3.13) $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}(\phi)+\Big\{2^{-n_{0}}\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}+T(1+\delta^{-1}+2^{n_{0}})\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}$
$+\big[T(1+\delta^{-1})+T^{2}(1+\delta^{-1})^{2}\big]\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}^{2}\Big\}$ $\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}.$
We now choose $n_{0}$ as $2^{n_{0}}=C(1+M_{0}^{2})$ with a large constant $C>0$ so that $2^{-n_{0}}\mathcal{E}^{2}(\phi)\leq 2^{-n_{0}}M_{0}^{2}\ll 1$, and choose $T_{0}$ sufficiently small so that $T_{0}(1+\delta^{-1})\big((1+2^{n_{0}})M_{0}^{2}+M_{0}^{4}\big)\ll 1$. Then, (3.13) and a standard continuity argument show that
$\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}\lesssim\mathcal{E}^{2}(\phi),\qquad 0<T<\min\{T_{0},T_{\ell}(\phi)\}.$
@@
The a priori bound in higher regularity (3.6) can be derived by a similar modified energy argument; we only outline the proof. It suffices to estimate the highest component $\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}$. Arguing similarly to the proof of (3.5), by Lemmas 2.1, 2.2 and interpolation inequalities, we deduce that
(3.14) $\frac{d}{dt}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\leq 2\beta\operatorname{Re}\big(\zeta_{\ell}(t)(iH)\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}+C_{k}(1+\delta^{-1})\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}.$
@@
$|\mathcal{I}_{k}(t)|\lesssim_{k}2^{-n_{1}}\mathcal{E}^{2}(u^{(\ell)}(t))\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},$
$\begin{aligned} \frac{d}{dt}\mathcal{I}_{k}(t)&=-2\beta\operatorname{Re}\big(iH\big(\zeta_{\ell}(t)\partial_{x}^{k+1}\overline{\zeta_{\ell}(t)}\big),\overline{\zeta_{\ell}(t)}\partial_{x}^{k}\zeta_{\ell}(t)\big)_{L^{2}}\\ &\quad+C_{k}\Big\{(1+2^{n_{1}})\mathcal{E}^{2}(u^{(\ell)}(t))+(1+\delta^{-1})\big[\mathcal{E}^{2}(u^{(\ell)}(t))\big]^{2}\Big\}\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}\end{aligned}$